You sell 310 pumpkins for $1950.00. Small pumpkins are $3.00 and large are $9.00. How many of each size did you sell
Answer by mananth(16949)

small pumpkin x numbers , Price = $3.00
large pumkin y numbers , Price = $9.00
Total 345
1 x + 1 y = 310 .............1
3 x + 9 y = 1950 .............2
Eliminate y
multiply (1)by -9
Multiply (2) by 1
-9 x -9 y = -2790
3 x 9 y = 1950
Add the two equations
-6 x = -840
/ -6
x = 140.00
plug value of x in (1)
1 x + 1 y = 310
140 + 1 y = 310
1 y = 310 -140
1 y = 170
y = 170.00
small pumpkin 140 numbers
large pumkin 170 numbers
m.ananth@hotmail.ca